the sum of x and 4x = 4. what is the value of x?
Answer by Cromlix(4381)

x + 4x = 4
5x = 4
Divide both sides by 5
x = 4/5
Check in original problem:-
x + 4x = 4
4/5 + 4(4/5) = 4
4/5 + 16/5 = 4
20/5 = 4
4 = 4